OpenStudy (anonymous):

WILL MEDAL AND FAN Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m2 by increasing the width and length by the same amount. What will be the length (longer dimension) of the new garden? Question 5 options: 16 m 17 m 18 m 19 m

OpenStudy (anonymous):

17 m

OpenStudy (anonymous):

the new dimensions of the garden will be (12+x)m wide and (14+x)m long Use the area formula A=lw (12+x)(14+x)=255m^2 x^2+26x-87=0 (x+29)(x-3)=0 x cannot be negative because she is adding to the dimensions so x=3. the new dimensions will be 15m wide x 17m long

OpenStudy (anonymous):

@Kyliejones youthere?

OpenStudy (anonymous):

Sorry but thank you!!